# Generalized saturation game

### v1

@@ 3.1 Counting $P_{3}$’s
The game graph at the end of the game is saturated and contains no $S_{4}$, so its maximum degree is at most $2$. Hence, at most one of its connected components is not a cycle – and it can be either an isolated edge or an isolated vertex. Counting $P_{3}$’s in such a graph with $n$ vertices, clearly, the number is either $n-2$, if there is an isolated edge, $n-1$, if there is an isolated vertex, or $n$, if all components are cycles.
To simplify the analysis of this game, we introduce some notation. We define the game deficit as $s(n,\#P_{3},S_{4})-n$. As we concluded above, the game deficit can be either $-2$, $-1$, or 00.
If we look at the game graph mid-game, the components that are cycles cannot be touched for the remainder of the game, and the number of $P_{3}$s in them is equal to the number of vertices. Therefore we can remove them from the graph for the rest of the game without influencing the deficit.
@@
For $n=4$ the deficit is 0. Max as first completes a $P_{4}$ in his second move, and Max as second claims an isolated edge in his first move. Then in both cases, the only saturated supergraph is $C_{4}$.
If $n=5$ and Max is first, he can complete a $P_{4}$ in his second move guaranteeing $[5]_{1}\geq-1$. On the other hand, Mini in her second move can complete a $P_{3}$ and then, if Max does not close it into a 3-cycle, Mini closes the long path (with less than $5$ vertices) into a cycle in her following move, guaranteeing $[5]_{1}\leq-1$.
For $n=5$ with Max second, he claims an isolated edge in his first move and then completes a $P_{5}$ in his second move, thus guaranteeing a zero deficit.
If $n=6$, Max as first makes sure that after his second move one of the two cases holds: either there is a triangle, in which case the game is reduced to $[3]_{2}$, or there are three isolated edges, which he can complete into a $P_{6}$ in his third move. In both cases there is a zero deficit.
When $n=6$ and Mini is first, she can make sure that she completes a $P_{4}$ in her second move. If Max does not close this path into a cycle there will be a long path after his move, so Mini can close the long path into a cycle (not spanning all the vertices), thus ensuring $[6]_{2}\leq-1$. On the other hand, Max can complete a $P_{3}$ in his first move, and unless Mini closes it into a triangle, Max can ensure a $P_{5}$ after his second move. Thus $[6]_{2}\geq-1$.
@@
If $n=7$ and Max is second, he can complete a $P_{3}$ in his first move. Then he can make sure that after his second move, the graph is either a triangle and an isolated edge, which by Observation 3.1 reduces to $[4]_{1}$, or it is a $C_{4}$, which reduces to $[3]_{2}$. ∎
@@
Figure 1: The Minimax game tree, giving a lower bound on the deficit for the three games in Lemma 3.3.
@@
From the beginning of the game, Max will follow the path extension strategy, making sure that after each of his moves the played edges form a number of cycles, possibly none, and exactly one path of positive length. Let us first show that, as long as there are at least two isolated vertices in the graph, he can follow his strategy. Indeed, independent of who starts the game, Max in his very first move can create a single path, either a $P_{2}$ or $P_{3}$. Then, if we assume that Max followed this strategy for a number of rounds, in her following move Mini can do one of three things: close a cycle, in which case Max claims an isolated edge, extend the path, in which case Max extends the path, or claim an isolated edge, in which case Max joins the isolated edge to the path, thus again extending it.
Max follows the path extension strategy until the number of isolated vertices drops below $7$. If $n$ and $i$ originally had different parity, then there are exactly six isolated vertices at that point, and the game is reduced to either $[6,1,0]_{2}$ or $[6,0,1]_{2}$. In $[6,1,0]_{2}$ Max can guarantee that after his move there is one $P_{4}$ and $4$ isolated vertices. Whatever Mini plays, Max can reduce the game to either $[3]_{2}$ or $[2,1]_{2}=[4]_{1}$, by Observation 3.1. In both cases the game deficit is zero. In $[6,0,1]_{2}$ Mini has several options. She can close the long path into a cycle, reducing the game to $[6]_{1}$. Alternatively, she can claim an isolated edge, to which Max can respond by closing the long path into a cycle, reducing the game to $[4,1,0]_{2}=[6]_{1}$, by Observation 3.1. Finally, she can add one edge to the long path, to which Max can again respond by closing the long path into a cycle, reducing the game to $[5]_{2}$. In all cases the game deficit is zero.
@@
If $n=2\ell$ and Mini is first to play, she makes sure to complete a $P_{4}$ after her second move. Then Max can play the following three moves. If he closes the path into a cycle, the game reduces to $[2\ell-4]_{2}$. If he claims an isolated edge, Mini closes the $P_{4}$ into a cycle, thus reducing the game to $[2\ell-6,1]_{1}=[2\ell-4]_{2}$, by Observation 3.1. Finally, if he extends the path into $P_{5}$, Mini closes it into a cycle reducing the game to $[2\ell-5]_{1}$.
If $n=2\ell+1$ and Mini is second to play, she completes a $P_{3}$ in her first move. Then similarly to the previous case, Max can play the following three moves. If he closes the path into a cycle, the game reduces to $[2\ell-2]_{2}$. If he claims an isolated edge, Mini closes the $P_{3}$ into a cycle, thus reducing the game to $[2\ell-4,1]_{1}=[2\ell-2]_{2}$. Finally, if he extends the path into $P_{4}$, Mini closes it into a cycle reducing the game to $[2\ell-3]_{1}$. ∎
@@ 3.2 Counting $P_{4}$’s
Let us first observe that in a saturated $S_{4}$-free graph the numbers of $P_{3}$’s and $P_{4}$’s are the same on all connected components except on triangles, where each such component has three $P_{3}$’s and zero $P_{4}$’s. One immediate consequence follows.
@@
On $4$ vertices, depending on who starts, Max can create either a $P_{4}$ or two disjoint edges, and in both cases, the only saturated supergraph is a $C_{4}$. On $5$ vertices, Max as second claims an isolated edge in his first move, and then creates a $P_{5}$ in his second move.
In the game $[8]_{1}$ Mini right away creates a $P_{3}$, and Max has to extend it to avoid the creation of a triangle. Then Mini further extends it, getting to three isolated vertices and a long path. From there, all possible moves result in a deficit of at least $-2$ as Mini will close the long path to a cycle.
In the game $[9]_{2}$, Max has to claim an isolated edge in his first move to avoid the creation of a triangle right away. If Mini then extends that path into a $P_{3}$, Max has to play adjacent to it. In any case Mini can complete a $P_{6}$ in her next move, and make sure that the long path is closed into a cycle in the round that follows. ∎
@@
If Max is second, he claims an isolated edge in his first move. Then after his second move, he can make sure that the played edges form either a $P_{5}$ or a $P_{4}+P_{2}$. In the former case, he applies the path extension strategy, and if Mini ever completes a cycle Max proceeds as first on the remaining isolated vertices. In the latter case, Mini can either create a $C_{4}$, or Max can respond by creating a $C_{4}+P_{2}+P_{2}$, or Max can respond by creating a $P_{7}$. All three cases are already analyzed above in Lemma 3.7 and all lead to a deficit of at least -3.∎
@@
Throughout the game, Mini extends the matching of the game graph by playing isolated edges. Note that to avoid the creation of a triangle, Max cannot connect an isolated edge with an isolated vertex. If he ever connects two isolated edges to create a $P_{4}$, Mini closes it into a $C_{4}$ and removes it from the game. Due to parity, towards the very end of the game there will be only one isolated edge and one isolated vertex, with Max to play, and the triangle will eventually be created. ∎
@@
Mini as second can create a $P_{3}$, and then Max has to extend it to $P_{4}$ to avoid the creation of a triangle. Mini then closes it into a $C_{4}$ and the game is reduced to $[n-4]_{1}$.
If Mini is first she can create a $P_{4}$ after two moves. Then if Max closes it into a $C_{4}$ or claims an isolated edge, Mini responds by doing the other one of those two things reducing the game to $[n-4]_{2}$. If, however, Max extends the $P_{4}$ into a $P_{5}$, Mini closes it into a cycle reducing the game to $[n-5]_{1}$. ∎
@@ 3.3 Counting $P_{5}$’s
    There is a $P_{i}\in S$, where $i\in\{3,4\}$. Mini ’closes’ such a path to a $C_{i}$.
@@
    $S$ consists of edges. Mini tries to claim an isolated $K_{2}$. If that is not possible, Mini connects one edge with an isolated vertex creating a $P_{3}$.
@@
    $S=\emptyset$. Mini plays an isolated $K_{2}$ if she can. If she can not, then the graph of the played edges consists of cycles and the game is over.
@@
    One $P_{3}$ and one $P_{4}$. Mini closes a $C_{4}$. Note that after Max’s next move, the game is continued in one of the following cases: 1, 4, or 5.
@@
    One $P_{3}$. Mini closes a triangle. Then until the end of the game, she just closes $C_{4}$’s.
@@
    One $P_{4}$. Mini closes it. Then until the end of the game, she just closes $C_{4}$’s.
@@
    One $P_{5}$. Mini closes a $C_{5}$. Note that after Max’s next move, the game is continued in: 3.
@@
    None. Mini connects two isolated edges and completes a $P_{4}$ and then until the end of the game, she always closes cycles. If there are two paths, she closes one arbitrarily.
@@
Observe that after Mini’s last move played in Stage 1 the graph of the played edges is a disjoint union of the following components, some $C_{3}$s, some $C_{4}$s, some $K_{2}$s, and at most one $P_{3}$. Since a cycle cannot be connected to any other component, all possible scenarios before the Mini’s first move in Stage 2 are listed in the five cases above.
If Mini applies one of the rules 2 or 3, it is clear that she can close the cycle and after that, only the pairs of isolated edges can be connected into one component, hence the only possible move for Max is to complete a $P_{4}$, and then Mini closes a $C_{4}$ and these moves will be repeated to the end of the game. Here the graph at the end of the game does not contain a $P_{5}$, hence the score is $s=0$.
If Mini applies rule 4, the score is $s=5$. If Mini applies rule 5, that is the only possible move for her and after that, Max can complete another $P_{4}$ component or create a $P_{6}$. In both of these cases, Mini will close a cycle in her following move. If she closes a $C_{6}$ (that is possible just once), the game is reduced to 3. Otherwise, Mini closes a $C_{4}$ and repeats closing a cycle such that at the end of the game the graph contains at most one $C_{6}$ and all the remaining cycles have at most $4$ vertices. Therefore, the score of the game is $s\leq 6$. This concludes the first part of the theorem.
@@
    If there is a $P_{i}$ component, where $i\geq 5$, Max closes a $C_{i}$ and the game is over.
@@
    Else, if Mini plays in a component that is a $P_{i}$ ($i\geq 3$) after her move, Max adds one isolated edge to this path, making a $P_{i+2}$, or if there is no isolated edge he adds an isolated vertex, making a $P_{i+1}$.
When Max has to apply rule 3 for the first time, Mini has completed either a $P_{3}$ or a $P_{4}$ component. In case this component is a $P_{3}$, there has to be at least one isolated edge because the first round has been finished with two isolated edges. Otherwise, if the component is a $P_{4}$ there must be either an isolated edge or vertex. Therefore, after this move of Max, there must be a $P_{i}$ component, where $i\geq 5$, so the game will be over in the next round with $s_{2}\geq 5$.
If the graph is matching and there is at most one isolated vertex, it must be Mini’s turn (because of the condition for $n$). Therefore, Mini has to complete either a $P_{3}$ or $P_{4}$, and in his following move, Max completes a $P_{i}$ where $i\geq 5$, following rule 3, so the game will be finished in the next round with $s_{2}\geq 5$. ∎

### v2

@@ 3.1 Counting $P_{3}$’s
The game graph at the end of the game is saturated and contains no copy of $S_{4}$, so its maximum degree is at most $2$. Hence, at most one of its connected components is not a cycle – and it can be either an isolated edge or an isolated vertex. Counting copies of $P_{3}$ in such a graph with $n$ vertices, clearly, the number is either $n-2$, if there is an isolated edge, $n-1$, if there is an isolated vertex, or $n$, if all components are cycles.
To simplify the analysis of this game, we introduce some notation. We define the game deficit as $s_{i}(n,\#P_{3},S_{4})-n$. As we concluded above, the game deficit can be either $-2$, $-1$, or 00.
If we look at the game graph mid-game, the components that are cycles cannot be touched for the remainder of the game, and the number of copies of $P_{3}$ in them is equal to the number of vertices. Therefore, we can remove them from the graph for the rest of the game without influencing the deficit.
@@
For $n=4$ the deficit is 0. Max as first completes a copy of $P_{4}$ in his second move, and Max as second claims an isolated edge in his first move. Then in both cases, the only saturated supergraph is $C_{4}$.
If $n=5$ and Max is first, he can complete a copy of $P_{4}$ in his second move guaranteeing $[5]_{1}\geq-1$. On the other hand, Mini in her second move can complete a copy of $P_{3}$ and then, if Max does not close it into a 3-cycle, Mini closes the long path (with less than $5$ vertices) into a cycle in her following move, guaranteeing $[5]_{1}\leq-1$.
For $n=5$ with Max second, he claims an isolated edge in his first move and then completes a copy of $P_{5}$ in his second move, thus guaranteeing a zero deficit.
If $n=6$, Max as first makes sure that after his second move one of the two cases holds: either there is a triangle, in which case the game is reduced to $[3]_{2}$, or there are three isolated edges, which he can complete into a copy of $P_{6}$ in his third move. In both cases there is a zero deficit.
When $n=6$ and Mini is first, she can make sure that she completes a copy of $P_{4}$ in her second move. If Max does not close this path into a cycle (otherwise the game is reduced to $[2]_{2}$), there will be a long path after his move, so Mini can close the long path into a cycle (not spanning all the vertices), thus ensuring $[6]_{2}\leq-1$. On the other hand, Max can complete a copy of $P_{3}$ in his first move, and unless Mini closes it into a triangle (in which case the game is reduced to $[3]_{1}$), Max can ensure a copy of $P_{5}$ after his second move. Thus $[6]_{2}\geq-1$.
@@
If $n=7$ and Max is second, he can complete a copy of $P_{3}$ in his first move. Then he can make sure that after his second move, the graph is either a triangle and an isolated edge, which by Observation 3.1 reduces to $[4]_{1}$, or it is a copy of $C_{4}$, which reduces to $[3]_{2}$. ∎
@@
Figure 1: The Minimax game tree, giving an upper bound on the deficit for the three games in Lemma 3.3.
@@
From the beginning of the game, Max will follow the path extension strategy, making sure that after each of his moves the played edges form a number of cycles, possibly none, and exactly one path of positive length. Let us first show that, as long as there are at least two isolated vertices in the graph, he can follow his strategy. Indeed, independent of who starts the game, Max in his very first move can create a single path, either a copy of $P_{2}$ or a copy of $P_{3}$. Then, if we assume that Max followed this strategy for a number of rounds, in her following move Mini can do one of three things: close a cycle, in which case Max claims an isolated edge, extend the path, in which case Max extends the path, or claim an isolated edge, in which case Max joins the isolated edge to the path, thus again extending it.
Max follows the path extension strategy until the number of isolated vertices drops below $7$. If $n$ and $i$ originally had different parity, then there are exactly six isolated vertices at that point, and the game is reduced to either $[6,1,0]_{2}$ or $[6,0,1]_{2}$. In $[6,1,0]_{2}$ Max can guarantee that after his move there is one copy of $P_{4}$ and $4$ isolated vertices. Whatever Mini plays, Max can reduce the game to either $[3]_{2}$ or $[2,1]_{2}=[4]_{1}$, by Observation 3.1. In both cases the game deficit is zero. In $[6,0,1]_{2}$ Mini has several options. She can close the long path into a cycle, reducing the game to $[6]_{1}$. Alternatively, she can claim an isolated edge, to which Max can respond by closing the long path into a cycle, reducing the game to $[4,1,0]_{2}=[6]_{1}$, by Observation 3.1. Finally, she can add one edge to the long path, to which Max can again respond by closing the long path into a cycle, reducing the game to $[5]_{2}$. In all cases the game deficit is zero.
@@
If $n=2\ell$ and Mini is first to play, she makes sure to complete a copy of $P_{4}$ after her second move. Then Max can play the following three moves. If he closes the path into a cycle, the game reduces to $[2\ell-4]_{2}$. If he claims an isolated edge, Mini closes the copy of $P_{4}$ into a cycle, thus reducing the game to $[2\ell-6,1]_{1}=[2\ell-4]_{2}$, by Observation 3.1. Finally, if he extends the path into $P_{5}$, Mini closes it into a cycle reducing the game to $[2\ell-5]_{1}$.
If $n=2\ell+1$ and Mini is second to play, she completes a copy of $P_{3}$ in her first move. Then similarly to the previous case, Max can play the following three moves. If he closes the path into a cycle, the game reduces to $[2\ell-2]_{2}$. If he claims an isolated edge, Mini closes the copy of $P_{3}$ into a cycle, thus reducing the game to $[2\ell-4,1]_{1}=[2\ell-2]_{2}$. Finally, if he extends the path into $P_{4}$, Mini closes it into a cycle reducing the game to $[2\ell-3]_{1}$. ∎
@@ 3.2 Counting $P_{4}$’s
Let us first observe that in a saturated $S_{4}$-free graph the numbers of copies of $P_{3}$ and $P_{4}$ are the same on all connected components except on triangles, where each such component has three copies of $P_{3}$ and zero copies of $P_{4}$. One immediate consequence follows.
@@
On $4$ vertices, depending on who starts, Max can create either a copy of $P_{4}$ or two disjoint edges, and in both cases, the only saturated supergraph is a copy of $C_{4}$.
On $5$ vertices, Max as second claims an isolated edge in his first move, and then creates a copy of $P_{5}$ in his second move. Max as first in his second move creates a copy of $P_{4}$ and thus ensures that the deficit is at least $-1$.
In the game $[8]_{1}$ Mini right away creates a copy of $P_{3}$, and Max has to extend it to avoid the creation of a triangle. Then Mini further extends it, getting to three isolated vertices and a long path. From there, all possible moves result in a deficit of at least $-2$ as Mini will close the long path to a cycle.
In the game $[9]_{2}$, Max has to claim an isolated edge in his first move to avoid the creation of a triangle right away. If Mini then extends that path into a copy of $P_{3}$, Max has to play adjacent to it. In any case Mini can complete a copy of $P_{6}$ in her next move, and make sure that the long path is closed into a cycle in the round that follows. ∎
@@
If Max is second, he claims an isolated edge in his first move. Then after his second move, he can make sure that the played edges form either a copy of $P_{5}$ or a $P_{4}+P_{2}$. In the former case, he applies the path extension strategy, and if Mini ever completes a cycle Max proceeds as first on the remaining isolated vertices. In the latter case, Mini can either create a copy of $C_{4}$, or Max can respond by creating a $C_{4}+P_{2}+P_{2}$, or Max can respond by creating a copy of $P_{7}$. If a copy of $P_{7}$ is created, Max simply follows the path extension strategy similarly as above, and if instead $C_{4}+P_{2}+P_{2}$ is created, one can ignore $C_{4}$ and hence proceed inductively back to the initial condition of this case.
@@
Throughout the game, Mini extends the matching of the game graph by playing isolated edges. Note that to avoid the creation of a triangle, Max cannot connect an isolated edge with an isolated vertex. If he ever connects two isolated edges to create a copy of $P_{4}$, Mini closes it into a copy of $C_{4}$ and removes it from the game. Due to parity, towards the very end of the game there will be only one isolated edge and one isolated vertex, with Max to play, and the triangle will eventually be created. ∎
@@
Mini as second can create a copy of $P_{3}$, and then Max has to extend it to a copy of $P_{4}$ to avoid the creation of a triangle. Mini then closes it into a copy of $C_{4}$ and the game is reduced to $[n-4]_{1}$.
If Mini is first she can create a copy of $P_{4}$ after two moves. Then if Max closes it into a $C_{4}$ or claims an isolated edge, Mini responds by doing the other one of those two things reducing the game to $[n-4]_{2}$. If, however, Max extends the copy of $P_{4}$ into a $P_{5}$, Mini closes it into a cycle reducing the game to $[n-5]_{1}$. ∎
@@ 3.3 Counting $P_{5}$’s
    $S$ contains a copy of $P_{i}$, where $i\in\{3,4\}$. Mini ’closes’ such a path to a copy of $C_{i}$.
@@
    $S$ consists of edges. Mini tries to claim an isolated $K_{2}$. If that is not possible, Mini connects one edge with an isolated vertex creating a copy of $P_{3}$.
@@
    $S$ is an empty set. Mini plays an isolated $K_{2}$ if she can. If she can not, then the graph of the played edges consists of cycles and the game is over.
@@
    One copy of $P_{3}$ and one copy of $P_{4}$. Mini closes a copy of $C_{4}$. Note that after Max’s next move, the game is continued in one of the following cases: 1, 4, or 5.
@@
    One copy of $P_{3}$. Mini closes a triangle. Then until the end of the game, she just closes copies of $C_{4}$.
@@
    One copy of $P_{4}$. Mini closes it. Then until the end of the game, she just closes copies of $C_{4}$.
@@
    One copy of $P_{5}$. Mini closes a copy of $C_{5}$. Note that after Max’s next move, the game is continued in: 3.
@@
    None. Mini connects two isolated edges and completes a copy of $P_{4}$ and then until the end of the game, she always closes cycles. If there are two paths, she closes one arbitrarily.
@@
Observe that after Mini’s last move played in Stage 1 the graph of the played edges is a disjoint union of the following components, some copies of $C_{3}$, some copies of $C_{4}$, some copies of $K_{2}$, and at most one copy of $P_{3}$. Since a cycle cannot be connected to any other component, all possible scenarios before the Mini’s first move in Stage 2 are listed in the five cases above.
If Mini applies one of the rules 2 or 3, it is clear that she can close the cycle and after that, only the pairs of isolated edges can be connected into one component, hence the only possible move for Max is to complete a copy of $P_{4}$, and then Mini closes a copy of $C_{4}$ and these moves will be repeated to the end of the game. Here the graph at the end of the game does not contain a copy of $P_{5}$, hence the score is $s=0$.
If Mini applies rule 4, the score is $s=5$. If Mini applies rule 5, that is the only possible move for her and after that, Max can complete another copy of $P_{4}$ component or create a copy of $P_{6}$. In both of these cases, Mini will close a cycle in her following move. If she closes a copy of $C_{6}$ (that is possible just once), the game is reduced to 3. Otherwise, Mini closes a copy of $C_{4}$ and repeats closing a cycle such that at the end of the game the graph contains at most one copy of $C_{6}$ and all the remaining cycles have at most $4$ vertices. Therefore, the score of the game is $s\leq 6$. This concludes the first part of the theorem.
@@
    If there is a $P_{i}$ component, where $i\geq 5$, Max closes a copy of $C_{i}$ and the game is over.
@@
    Else, if Mini plays in a component that is a copy of $P_{i}$ ($i\geq 3$) after her move, Max adds one isolated edge to this path, making a copy of $P_{i+2}$, or if there is no isolated edge he adds an isolated vertex, making a copy of $P_{i+1}$.
When Max has to apply rule 3 for the first time, Mini has completed either a copy of $P_{3}$ or a copy of $P_{4}$. In case this component is a copy of $P_{3}$, there has to be at least one isolated edge because the first round has been finished with two isolated edges. Otherwise, if the component is a copy of $P_{4}$ there must be either an isolated edge or vertex. Therefore, after this move of Max, there must be a $P_{i}$ component, where $i\geq 5$, so the game will be over in the next round with $s_{2}\geq 5$.
If the graph is matching and there is at most one isolated vertex, it must be Mini’s turn (because of the condition for $n$). Therefore, Mini has to complete either a copy of $P_{3}$ or a copy of $P_{4}$, and in his following move, Max completes a copy of $P_{i}$ where $i\geq 5$, following rule 3, so the game will be finished in the next round with $s_{2}\geq 5$. ∎